OpenStudy (anonymous):

Show how you solve the following two equations:

OpenStudy (anonymous):

?????

OpenStudy (anonymous):

these are 2 different ones here x-15=32 x-7.2=23.1

OpenStudy (anonymous):

I don't remember how to do it

OpenStudy (anonymous):

Simplifying X + -15 = 32 Reorder the terms: -15 + X = 32 Solving -15 + X = 32 Solving for variable 'X'. Move all terms containing X to the left, all other terms to the right. Add '15' to each side of the equation. -15 + 15 + X = 32 + 15 Combine like terms: -15 + 15 = 0 0 + X = 32 + 15 X = 32 + 15 Combine like terms: 32 + 15 = 47 X = 47 Simplifying X = 47

OpenStudy (anonymous):

Simplifying X + -15 = 32 Reorder the terms: -15 + X = 32 Solving -15 + X = 32 Solving for variable 'X'. Move all terms containing X to the left, all other terms to the right. Add '15' to each side of the equation. -15 + 15 + X = 32 + 15 Combine like terms: -15 + 15 = 0 0 + X = 32 + 15 X = 32 + 15 Combine like terms: 32 + 15 = 47 X = 47 Simplifying X = 47

OpenStudy (anonymous):

give me a medal ???